OpenStudy (anonymous):

Calculate the wavelength, λ, in meters of a photon capable of exciting an electron in He+ from the ground state to n = 4.

OpenStudy (anonymous):

I have used the equation RZ^2 (1/ni^2-1/nf^2)=4.113*10^7 1/4.113*10^7=2.43*10^-6 which is not correct..can someone please let me know what I did wrong? thank you

OpenStudy (aaronq):

use the rydberg equation,

OpenStudy (aaronq):

someone else posted a similar question

OpenStudy (anonymous):

I know...thats what I used...and it s still incorrect hence I thought that my strategy might be wrong...

OpenStudy (aaronq):

the final and initial placements are incorrect, it's (1/nf^2 - 1/ ni^2) so final minus initial

OpenStudy (anonymous):

thank you so very much...

OpenStudy (aaronq):

no prob!